# Regularizing Effect for a Nonlocal Maxwell-Schrödinger System

### v1

@@ 1 Introduction
In this paper, we investigate the existence and regularity of positive solutions for the following class of Maxwell-Schrödinger systems of Kirchhoff
@@
where, from now on, $\Omega\subset\mathbb{R}^{N}$, is an open bounded subset, $N>2$, $1<p<N$ and $f\in L^{m}(\Omega)$ for $m>1$. Moreover, $g$ and $h:\Omega\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ are both Carathéodory functions, which may be non monotone, responsible for the coupling of our system, see the assumptions below.
In recent years, there have been several results investigating the gain of regularity of solutions of Maxwell-Schrödinger systems. As it turned out, even in rough regimes where the datum is not regular, the coupling in these systems allows the existence of solutions which somehow are more regular than expected, the so–called Regularizing Effect. By now, it is known that this is caused by the very interesting nature of the coupling between the equations of systems of Maxwell-Schrödinger type. Without the intention of being complete, the interested reader is invited to see [1, 3, 6, 8, 9, 10, 13] plus the references therein, and many others.
The starting point for this analysis is due to L. Boccardo in [3], where this phenomenon was first investigated. Later, in a series of works in collaboration with L. Orsina, other aspects of regularity gains were explored, for instance see [6, 7], or more recently [9], and the references therein. Summarizing, the authors prove existence/ regularizing properties of/for solutions of local systems of Maxwell-Schrödinger type. In order to illustrate the ideas, we quote that in [6], they address the following problem
@@
where, from now on, $M(x)$ is a uniformly elliptic symmetric bounded matrix, i.e., $M(x)\xi\cdot\xi\geqslant\alpha|\xi|^{2}\mbox{ and }|M(x)|\leqslant\beta\ \ \mbox{a.e in}\ \Omega\ \mbox{and}\ \forall\ \xi\in\mathbb{R}^{N}$, for $0<\alpha\leqslant\beta$, $A$ is a positive constant, and $f\in L^{m}(\Omega)$. In short, among other results, the authors prove that even when $m<(2^{*})^{\prime}$, if $r>2^{*}$ and $r^{\prime}\leqslant m$, there exist $u,\varphi\in W^{1,2}_{0}(\Omega)$, weak solution of (1.1). This is interesting since $f$ lies below the dual threshold $(2^{*})^{\prime}$, $|u|^{r}$ is a priori low summable, and also because the weight $A\varphi$ is not assumed to be regular or strictly positive. Moreover, when $1<r<2^{*}$, if $\max\{1,\frac{Nr}{N+2r}\}<m<(2^{*})^{\prime}$ there exists $(u,v)\in W^{1,m^{*}}_{0}(\Omega)\times W^{1,q}_{0}(\Omega)$, solution of (1.1) for
@@
In the particular case where $m\geqslant(r+\theta+1)^{\prime}$, $0<\theta<\min\big{(}\frac{4}{N-2},1\big{)}$, $12^{*}<r+\theta+$ 1 and $(r+\theta+1)^{\prime}\leqslant m<(2^{*})^{\prime}$, recalling that $f\in L^{m}(\Omega)$, the authors prove the existence of $(u,v)\in W^{1,2}_{0}(\Omega)\times W^{1,2}_{0}(\Omega)$, a regularized weak solution of (1.2).
For further developments on the local case, see [1], [7], and [10]. In [1], in cooperation with D. Arcoya, by assuming a interesting growth condition on $f(x)$, it is addressed a version of (1.1) with the constant $A>0$ replaced by a $A(x)\in L^{1}(\Omega)$, $A(x)\geqslant 0$, and also with a modification of the diffusion term on the second equation. In [7], the authors include a sublinear and singular structures in (1.1), i.e., the nonlinearities become $A\varphi^{\theta}u^{r-1}$ and $\frac{u^{r}}{\varphi^{1-\theta}}$ where $\theta\in(0,1).$ Moreover, for an investigation of a quasilinear version of (1.1), see [10].
@@
$\begin{cases}&-\mbox{div}\Big{(}\big{[}a(x)+\|\nabla u\|^{2}_{L^{2}}\big{]}\nabla u\Big{)}+\varphi|u|^{r-2}u=f\ \ \ \ \ \mbox{in}\ \ \ \Omega;\\ &-\mbox{div}(M(x)\nabla\varphi)=|u|^{r}\ \ \ \ \ \ \ \mbox{in}\ \ \ \ \ \Omega;\\ &u=v=0\ \ \ \ \ \ \mbox{on}\ \ \ \ \ \partial\Omega,\end{cases}$ (1.3)
a Kirchhoff–Schrödinger–Maxwell system, where $a:\Omega\to\mathbb{R}$ is a measurable function such that $0<\alpha\leqslant a(x)\leqslant\beta.\ \ x$ a.e in $\Omega$, see (1.1) above. The authors conduct a through analysis of existence, nonexistence and regularity of solutions of (1.3) for $1<m<\infty$ and $r>1$. In particular, regarding regularizing zones, among other results, the authors proved that, if $(r+1)^{\prime}\leqslant m<(2^{*})^{\prime}$ and $r\geqslant 2^{*}-1$, or if $(2^{*})^{\prime}\leqslant m$ and $11<r<2^{*}-$ 1, then there exists $(u,\varphi)\in W^{1,2}_{0}(\Omega)\times W^{1,2}_{0}(\Omega)$, solution of (1.3), where $u\in L^{\sigma}(\Omega)$ with $\sigma=\max\{{m^{**},\frac{m(2r+1)}{m+1}}\}$. In the case $m<(2^{*})^{\prime}$, clearly $\sigma=\frac{m(2r+1)}{m+1}$ ensuring a double regularizing effect: there exist energy solutions despite that $f\notin W^{-1,2}(\Omega)$ and the $u$ has improved integrability since $\sigma>m^{**}$, see Definition 1.2.
Further, let us add that as pointed out in [8], outside the energy zone, e.g., $1<m<\min\{(r+1)^{\prime},(2^{*})^{\prime}\}$, there might be interesting convergence phenomenon. Indeed, let us consider for $f\in L^{1}(\Omega)\setminus W^{-1,2}(\Omega)$, $21<r<$ 2 and $N=6$. In this case, they construct a sequence of approximate solutions $(u_{n},\varphi_{n})$ where $u_{n},\varphi_{n}\in W^{1,2}_{0}(\Omega)\cap L^{\infty}(\Omega)$. As it turned out, under these assumptions, they prove that $\|\nabla u_{n}\|^{2}_{L^{2}}\to+\infty$ and $u_{n}\rightharpoonup 0$ weakly in $W^{1,q}_{0}(\Omega)$, for $1<q<\frac{18}{11}$. Yet, by setting $v_{n}=\|\nabla u_{n}\|^{2}_{L^{2}}u_{n}$, they show that $v_{n}\rightharpoonup w$, the entropy solution of $-\Delta w=f$ and moreover, since $|u_{n}|^{r}\to 0$ strongly in $L^{1}(\Omega)$, then $\varphi_{n}\rightharpoonup 0$ weakly in $W^{1,\rho}_{0}(\Omega)$, $1<\rho<\frac{6}{5}$. As a consequence, the second equation degenerates and converges to $0=0$.
In the present paper, we are interested in the analysis of regularity properties of solutions to a quasilinear version of the nonlocal Maxwell-Schrodinger system given by (P). As a result, we have to deal with certain obstacles arising from the modifications that we introduced.If on one hand, we combine ideas from [8, 10] and [13], at the same time, we have to circumvent that our operators are not self-adjunct, providing some challenges when translating duality arguments from Laplacian to $p$-Laplacian equations, and yet work with a class of zeroth-order of nonlinearities. In particular, this makes the approach based on approximation and passage to the limit slightly more complex.
@@
We consider $r>1$, $10<\theta<p-$ 1, $1<p<N$, $1<m<\frac{N}{P}$, $g:\Omega\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $h:\Omega\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, both Carathéodory, satisfying
@@
Let us briefly remark that the prototypes for $g$ and $h$ are given by $g(s,t)=s|s|^{r-2}|t|^{\theta+1}$ and $h(s,t)=t|s|^{r}|t|^{\theta-1}$. Moreover, we stress that ($\mbox{H}_{1}$)-($\mbox{H}_{4}$) are slightly weaker than the hypotheses assumed in [13], see ($\mbox{H}^{\prime}_{1}$)-($\mbox{H}^{\prime}_{4}$). Grossly speaking, we now allow non-monotone nonlinearities which may change sign. For example, we can consider
@@
Let $1<p<N$. We say that $(u,v)$ in $W^{1,p}_{0}(\Omega)\times W^{1,p}_{0}(\Omega)$, is a solution for problem (P) if and only if
$\begin{cases}&\int_{\Omega}\bigg{(}\|\nabla u\|^{p}_{L^{p}}+\|\nabla v\|^{p}_{L^{p}}\bigg{)}|\nabla u|^{p-2}\nabla u\cdot\nabla\varphi+\int_{\Omega}g(x,u,v)\varphi=\int_{\Omega}f\varphi\ \ \ \ \forall\ \varphi\in C^{\infty}_{c}(\Omega)\\ &\int_{\Omega}\bigg{(}\|\nabla u\|^{p}_{L^{p}}+\|\nabla v\|^{p}_{L^{p}}\bigg{)}|\nabla v|^{p-2}\nabla v\cdot\nabla\psi=\int_{\Omega}h(x,u,v)\psi\ \ \ \ \ \forall\ \psi\in C^{\infty}_{c}(\Omega).\end{cases}$ ($P_{F}$)
where $g(.,u,v)$ and $h(.,u,v)\in L^{1}_{loc}(\Omega)$.
Remark that the solutions defined above is intermediate. Indeed, on one hand they are not regular enough to be considered as weak solutions, for instance we cannot take test functions in $W^{1,p}_{0}(\Omega)$. On the other hand, they are better than standard distributional or even “very weak” solutions since $(u,v)\in W^{1,p}_{0}(\Omega)\times W^{1,p}_{0}(\Omega)$.
@@
$\tau_{p}^{*}=\tau^{*}(p-1)=\dfrac{N\tau(p-1)}{N-\tau}\mbox{ and }\tau^{**}_{p}=(\tau_{p}^{*})^{*}=\dfrac{N\tau(p-1)}{N-\tau p}.$
@@
Consider $w\in W^{1,1}_{0}(\Omega)$ a distribution solution of
@@
Let $f\in L^{m}(\Omega),$ where $f\geqslant 0$ a.e in $\Omega,$ $\min\{(r+\theta+1)^{\prime},(p^{*})^{\prime}\}<m<\frac{N}{p}$ and $0<\theta<\min\{p-1,\frac{p^{2}}{N-p}\}.$ Then there exists a distributional solution $(u,v)$ for (P), with $u\in W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega)$, $u\geqslant 0$ a.e in $\Omega$ and $v\in W^{1,p}_{0}(\Omega)$, $v\geqslant 0$ a.e. in $\Omega$.
Our approach to prove Theorem 1.3 is based on the analysis of approximate problems by means of integral estimates obtained by the careful choice of test functions. Of course, in the end, we pass to the limit.
This theorem connects results from [8], [10] and [13], since it reaches a nonlocal, $p$-Laplacian-like Maxwell-Schrödinger system, for a class of nonlinearities. This combination, brings new difficulties to a problem which is already complex. If on one hand, we consider $g$ and $h$ satisfying weaker conditions than previous results, on the other hand, the coupling between these nonlinearities and the nonlocal terms makes the problem even more challenging, especially the passage to the limit of the approximate problems, which is very tricky. Regarding the restriction on $m>m_{0}=(r+\theta+1)^{\prime}$ mentioned before, this is considered to avoid known degeneration issues for limit solutions of Kirchhoff-Maxwell-Schrödinger systems, as explained above and described in [8].
@@ 2 Preliminaries
Further, consider $\varphi,\psi\in W^{1,p}_{0}(\Omega)$, such that $\|\varphi\|_{W^{1,p}_{0}}+\|\psi\|_{W^{1,p}_{0}}\leq 1$, $i=1,2$ and $u,v\in W^{1,p}_{0}(\Omega)$, where $\|u\|_{W^{1,p}_{0}}+\|v\|_{W^{1,p}_{0}}\leqslant\rho$, for a given $\rho>0$. Thus, by recalling that $|g_{\eta}(x,u,v)|\leqslant\eta$ and $|h_{\eta}(x,u,v)|\leqslant\eta$, a.e. in $\Omega$, by standard applications of Hölder’s and Sobolev’s inequalities, we arrive at
@@
Hence, $T$ is pseudomonotone and by the Minty–Browder Theorem, there exists $(u,v)\in W^{1,p}_{0}(\Omega)\times W^{1,p}_{0}(\Omega)$ satisfying (2.1). Moreover, since $F\in L^{t}(\Omega)$ where $t\geqslant(p^{*})^{\prime}$, by taking $\varphi=u$ in (2.1) and $\psi=v$ in (2.2), so that by ($\mbox{H}_{1}$), ($\mbox{H}_{4}$), and straightforward manipulations we have
@@
while $u\in L^{\infty}(\Omega)$ together with $\mbox{H}_{4}$ guarantees that $H(x,s)\leqslant d_{2}\tau\|u\|_{L^{\infty}}^{r}|s|^{\theta}+1$ a.e. in $\Omega$, which is uniform with respect to $\eta$.
@@
$\begin{cases}\ -\mbox{div}\bigg{(}\mathcal{A}_{k}(u_{k},v_{k})|\nabla u_{k}|^{p-2}\nabla u_{k}\bigg{)}+g(x,u_{k},v_{k})=f_{k}\mbox{ in }\Omega,\\ \ -\mbox{div}\bigg{(}\mathcal{A}_{k}(u_{k},v_{k})|\nabla v_{k}|^{p-2}\nabla v_{k}\bigg{)}=h(x,u_{k},v_{k})+\frac{1}{k}\mbox{ in }\Omega,\\ u_{k}=v_{k}=0\mbox{ on }\partial\Omega,\end{cases}$ (2.6)
@@
As a first step, we prove existence. For this, let us start by pointing out that $f_{k}\in L^{\infty}(\Omega)$ and that $|f_{k}|\leqslant|f|$ a.e. in $\Omega$, for all $k>0$, since $T_{k}(s)=\max(-k,\min(s,k))$.
@@
where $C$ is the maximum between the constants in (2.3), (2.4), $c_{2}$ and $d_{2}$. Hence, by combining Proposition 2.1, there exist $u_{k}$, $v_{k}\in W^{1,p}_{0}(\Omega)\cap L^{\infty}(\Omega)$ solution of (2.1) and (2.2). Moreover, it is clear that $g_{\tau}(.,u_{k},v_{k})=g(.,u_{k},v_{k})$ and $h_{\tau}(.,u_{k},v_{k})=h(.,u_{k},v_{k})$ a.e. in $\Omega$. Indeed, by recalling $\mbox{H}_{2}$, $\mbox{H}_{4}$, (2.3) and (2.4) we get $|g(x,u_{k},v_{k})|\leqslant c_{2}|u_{k}|^{r-1}|v_{k}|^{\theta+1}\leqslant C(k^{q}+1)<\eta$ and $|h(x,u_{k},v_{k})|\leqslant d_{2}|u_{k}|^{r}|v_{k}|^{\theta}\leqslant C(k^{q}+1)<\eta$ a.e. in $\Omega$, proving the claim. In this way, $(u_{k},v_{k})$ is a weak solution of (2.6).

### v2

# Regularizing Effect for a Nonlocal Maxwell–Schrödinger System

@@ 1 Introduction
In this paper, we investigate the existence and regularity of positive solutions for the following class of Maxwell-Schrödinger systems of Kirchhoff type
@@
[…]
In recent years, there have been several results investigating the gain of regularity of solutions of Maxwell–Schrödinger systems. As it turned out, even in rough regimes where the datum is not regular, the coupling in these systems allows the existence of solutions which somehow are more regular than expected, the so–called Regularizing Effect. By now, it is known that this is caused by the very interesting nature of the coupling between the equations of systems of Maxwell–Schrödinger type. Without the intention of being complete, the interested reader is invited to see [1, 3, 6, 8, 9, 10, 13] plus the references therein, and many others.
The starting point for this analysis is due to L. Boccardo in [3], where this phenomenon was first investigated. Later, in a series of works in collaboration with L. Orsina, other aspects of regularity gains were explored, for instance see [6, 7], or more recently [9], and the references therein. Summarizing, the authors prove existence/ regularizing properties of/for solutions for a class of local Maxwell–Schrödinger systems. In order to illustrate the ideas, we quote that in [6], they address the following problem
@@
where, from now on, $M(x)$ is a uniformly elliptic symmetric bounded matrix, i.e., $M(x)\xi\cdot\xi\geqslant\alpha|\xi|^{2}\mbox{ and }|M(x)|\leqslant\beta\ \ \mbox{a.e in}\ \Omega\ \mbox{and}\ \forall\ \xi\in\mathbb{R}^{N}$, for $0<\alpha\leqslant\beta$, $A$ is a positive constant, and $f\in L^{m}(\Omega)$. In short, among other results, the authors prove that even when $m<(2^{*})^{\prime}$, if $r>2^{*}$ and $r^{\prime}\leqslant m$, there exist $u,\varphi\in W^{1,2}_{0}(\Omega)$, weak solution of (1.1). This is interesting since $f$ lies below the dual threshold $(2^{*})^{\prime}$, $|u|^{r}$ is a priori low summable, and also because the weight $A\varphi$ is not assumed to be regular or strictly positive. Moreover, when $1<r<2^{*}$, if $\max\{1,\frac{Nr}{N+2r}\}<m<(2^{*})^{\prime}$ there exists $(u,\varphi)\in W^{1,m^{*}}_{0}(\Omega)\times W^{1,q}_{0}(\Omega)$, solution of (1.1) for
@@
In the particular case where $m\geqslant(r+\theta+1)^{\prime}$, $0<\theta<\min\big{(}\frac{4}{N-2},1\big{)}$, $2^{*}<r+\theta+1$ and $(r+\theta+1)^{\prime}\leqslant m<(2^{*})^{\prime}$, recalling that $f\in L^{m}(\Omega)$, the authors prove the existence of $(u,v)\in W^{1,2}_{0}(\Omega)\times W^{1,2}_{0}(\Omega)$, a regularized weak solution of (1.2).
For further developments on the local case, see [1], [7], and [10]. In [1], in cooperation with D. Arcoya, by assuming a interesting growth condition on $f(x)$, it is addressed a version of (1.1) with the constant $A>0$ replaced by a $A(x)\in L^{1}(\Omega)$, $A(x)\geqslant 0$, and also with a modification of the diffusion term on the second equation. In [7], the authors include sublinear and singular structures in (1.1), i.e., the nonlinearities become $A\varphi^{\theta}u^{r-1}$ and $\frac{u^{r}}{\varphi^{1-\theta}}$ where $\theta\in(0,1).$ Moreover, for an investigation of a quasilinear version of (1.1), see [10].
@@
$\begin{cases}&-\mbox{div}\Big{(}\big{(}a(x)+\|\nabla u\|^{2}_{L^{2}}\big{)}\nabla u\Big{)}+\varphi|u|^{r-2}u=f\ \ \ \ \ \mbox{in}\ \ \ \Omega;\\ &-\mbox{div}(M(x)\nabla\varphi)=|u|^{r}\ \ \ \ \ \ \ \mbox{in}\ \ \ \ \ \Omega;\\ &u=v=0\ \ \ \ \ \ \mbox{on}\ \ \ \ \ \partial\Omega,\end{cases}$ (1.3)
a Kirchhoff–Schrödinger–Maxwell (KMS) system, where $a:\Omega\to\mathbb{R}$ is a measurable function such that $0<\alpha\leqslant a(x)\leqslant\beta.\ \ x$ a.e in $\Omega$, see (1.1) above. The authors conduct a through analysis of existence, nonexistence and regularity of solutions of (1.3) for $1<m<\infty$ and $r>1$. In particular, regarding regularizing zones, among other results, the authors proved that, if $(r+1)^{\prime}\leqslant m<(2^{*})^{\prime}$ and $r\geqslant 2^{*}-1$, or if $(2^{*})^{\prime}\leqslant m$ and $1<r<2^{*}-1$, then there exists $(u,\varphi)\in W^{1,2}_{0}(\Omega)\times W^{1,2}_{0}(\Omega)$, solution of (1.3), where $u\in L^{q}(\Omega)$ for $q=\max\{{m^{**},\frac{m(2r+1)}{m+1}}\}$. In the case $m<(2^{*})^{\prime}$, clearly $q=\frac{m(2r+1)}{m+1}$ ensuring a double regularizing effect: there exist energy solutions despite that $f\notin W^{-1,2}(\Omega)$ and the $u$ has improved integrability since $q>m^{**}$, see Definition 1.2.
[…]
In the present paper, we are interested in the analysis of regularity properties of solutions to a quasilinear version of the nonlocal Maxwell–Schrodinger system given by (P). As a result, we have to deal with certain obstacles arising from the modifications that we introduced.If on one hand, we combine ideas from [8, 10] and [13], at the same time, we have to circumvent that our operators are not self–adjunct, causing some challenges when translating duality arguments from Laplacian to $p$-Laplacian equations, and yet work with a class of zeroth-order of nonlinearities. In particular, this makes the approach based on approximation and passage to the limit slightly more complex.
@@
[…]
@@
Let us briefly remark that the prototypes for $g$ and $h$ are given by $g(s,t)=s|s|^{r-2}|t|^{\theta+1}$ and $h(s,t)=t|s|^{r}|t|^{\theta-1}$. Moreover, we stress that ($\mbox{H}_{1}$)-($\mbox{H}_{4}$) are slightly weaker than the hypotheses assumed in [13], see ($\mbox{H}^{\prime}_{1}$)-($\mbox{H}^{\prime}_{4}$). Grossly speaking, we now allow nonmonotone nonlinearities which may change sign. For example, we can consider
@@
[…]
$\displaystyle\begin{cases}&\displaystyle\mathcal{A}(u,v)\int_{\Omega}|\nabla u|^{p-2}\nabla u\cdot\nabla\varphi+\int_{\Omega}g(x,u,v)\varphi=\int_{\Omega}f\varphi\ \ \ \ \forall\ \varphi\in C^{\infty}_{c}(\Omega)\\ &\displaystyle\mathcal{A}(u,v)\int_{\Omega}|\nabla v|^{p-2}\nabla v\cdot\nabla\psi=\int_{\Omega}h(x,u,v)\psi\ \ \ \ \ \forall\ \psi\in C^{\infty}_{c}(\Omega),\end{cases}$ ($P_{F}$)
$g(.,u,v)$ and $h(.,u,v)\in L^{1}_{loc}(\Omega)$, where $\mathcal{A}(u,v)=\|\nabla u\|^{p}_{L^{p}}+\|\nabla v\|^{p}_{L^{p}}$.
Remark that the latter definition of solutions is intermediate. Indeed, on one hand they are not regular enough to be considered as weak solutions, for instance, we cannot take test functions in $W^{1,p}_{0}(\Omega)$. On the other hand, they are better than standard distributional or even “very weak” solutions since $(u,v)\in W^{1,p}_{0}(\Omega)\times W^{1,p}_{0}(\Omega)$.
@@
$\tau_{p}^{*}=\tau^{*}(p-1)=\dfrac{N\tau(p-1)}{N-\tau}\mbox{ and }\tau^{**}_{p}=(\tau_{p}^{*})^{*}=\dfrac{N\tau(p-1)}{N-\tau p}.$ (1.4)
@@
Consider $w\in W^{1,1}_{0}(\Omega)$ a distributional solution of
@@
[…]
Our approach to prove Theorem 1.3 is based on the analysis of approximate problems allowing us to gather delicate integral estimates. This is achieved by means of the careful choice of tailored test functions. Of course, in the end, we pass to the limit.
This theorem connects results from [8], [10] and [13], since it reaches a nonlocal, $p$-Laplacian-like Maxwell-Schrödinger system, for a broader class of nonlinearities. This combination, brings new difficulties to a problem which is already complex. If on one hand, we consider $g$ and $h$ satisfying weaker conditions than previous results, on the other hand, the coupling between these nonlinearities and the nonlocal terms makes the problem even more challenging, especially the passage to the limit of the approximate problems, which is very tricky. Regarding the restriction on $m>m_{0}=(r+\theta+1)^{\prime}$ mentioned before, this is considered to avoid known degeneration issues for limit solutions of Kirchhoff-Maxwell-Schrödinger systems, as explained above and described in [8].
@@ 2 Preliminaries
Further, consider $\varphi,\psi\in W^{1,p}_{0}(\Omega)$, such that $\|\varphi\|_{W^{1,p}_{0}}+\|\psi\|_{W^{1,p}_{0}}\leq 1$, and $u,v\in W^{1,p}_{0}(\Omega)$, where $\|u\|_{W^{1,p}_{0}}+\|v\|_{W^{1,p}_{0}}\leqslant\rho$, for a given $\rho>0$. Thus, by recalling that $|g_{\eta}(x,u,v)|\leqslant\eta$ and $|h_{\eta}(x,u,v)|\leqslant\eta$, a.e. in $\Omega$, by standard applications of Hölder’s and Sobolev’s inequalities, we arrive at
@@
Hence, $T$ is pseudomonotone and by the Minty–Browder Theorem, there exists $(u,v)\in W^{1,p}_{0}(\Omega)\times W^{1,p}_{0}(\Omega)$ satisfying (2.1)-(2.2). Moreover, since $F\in L^{t}(\Omega)$ where $t\geqslant(p^{*})^{\prime}$, by taking $\varphi=u$ in (2.1) and $\psi=v$ in (2.2), so that by ($\mbox{H}_{1}$), ($\mbox{H}_{4}$), and straightforward manipulations we have
@@
while $u\in L^{\infty}(\Omega)$ together with ($\mbox{H}_{4}$) guarantees that $H(x,s)\leqslant d_{2}\tau\|u\|_{L^{\infty}}^{r}|s|^{\theta}+1$ a.e. in $\Omega$, which is uniform with respect to $\eta$.
@@
$\begin{cases}\ -\mbox{div}\Big{(}\mathcal{A}_{k}(u_{k},v_{k})|\nabla u_{k}|^{p-2}\nabla u_{k}\Big{)}+g(x,u_{k},v_{k})=f_{k}\mbox{ in }\Omega,\\ \ -\mbox{div}\Big{(}\mathcal{A}_{k}(u_{k},v_{k})|\nabla v_{k}|^{p-2}\nabla v_{k}\Big{)}=h(x,u_{k},v_{k})+\frac{1}{k}\mbox{ in }\Omega,\\ u_{k}=v_{k}=0\mbox{ on }\partial\Omega,\end{cases}$ (2.6)
@@
As a first step, we prove existence. For this, let us start by pointing out that $f_{k}\in L^{\infty}(\Omega)$ and that $|f_{k}|\leqslant\min\{|f|,k\}$ a.e. in $\Omega$, for all $k>0$, since $T_{k}(s)=\max(-k,\min(s,k))$.
@@
where $C$ is the maximum between the constants in (2.3), (2.4), $c_{2}$ and $d_{2}$. Hence, by Proposition 2.1, there exist $u_{k}$, $v_{k}\in W^{1,p}_{0}(\Omega)\cap L^{\infty}(\Omega)$ solution of (2.1) and (2.2). Moreover, it is clear that $g_{\tau}(.,u_{k},v_{k})=g(.,u_{k},v_{k})$ and $h_{\tau}(.,u_{k},v_{k})=h(.,u_{k},v_{k})$ a.e. in $\Omega$. Indeed, by recalling ($\mbox{H}_{2}$), ($\mbox{H}_{4}$), (2.3) and (2.4) we get $|g(x,u_{k},v_{k})|\leqslant c_{2}|u_{k}|^{r-1}|v_{k}|^{\theta+1}\leqslant C(k^{q}+1)<\eta$ and $|h(x,u_{k},v_{k})|\leqslant d_{2}|u_{k}|^{r}|v_{k}|^{\theta}\leqslant C(k^{q}+1)<\eta$ a.e. in $\Omega$, proving the claim. In this way, $(u_{k},v_{k})$ is a weak solution of (2.6).